From GlueXWiki
Jump to: navigation, search

July 16, 2015 Drift Chamber meeting


  1. Instructions for Bluejeans meeting connection
  2. Meeting ID: 290664653
  3. To join via a Web Browser, go to the page [1] https://bluejeans.com/290664653.


  1. CDC analysis update (Mike, Simon)
    1. cdc resolution plots (Simon)
  2. FDC analysis update (Luke,Lubomir)
  3. Electronics (Fernando, Chris, Nick)
  4. Engineering
  5. Other


Participants: Curtis, Mike(CMU), Sean(NU), Fernando, Luke, Dave, Simon, Beni, and Lubomir (JLab).

CDC update

- Simon presented latest results on the CDC resolution as a function of the distance to the wire. The log entry is linked above. He has an option in the analysis code to exclude one ring at a time from the tracking, while still calculating the residuals for the tubes in that ring. The resolutions are at/below 200 microns except for hits close to the wire. Improvement compared to previous results - procedure for smoothing the reconstructed track.

- Mike showed once more his slides presented at the last calibration meeting about the evidence for possible tube deformation. As pointed out before by Naomi, the drift time spectra on the different tubes are different, pointing to a possible tube deformations. There are "good" tubes with spectra as expected by Garfield, and "bad" tubes with long tails going beyond maximal expected drift time. Mike looked further at the azimuthal distribution of the hits vs drift time. Certainly there's a difference between the good and bad tubes. Much more statistics is needed to figure out the azimuthal and z-dependence of the deformations.

FDC update

- Luke performed measurements with 55FE source on the spare package with 60/40 Ar/CO2 for different HV. He described the calibration procedure and showed some results for 2050V and 1950V. The up/downstream avalanche peak position are well visible as with 40/60 gas mixture. At 2050V they are ~150microns apart (compared to 250 microns with 40/60 at almost the same gas gain), while at 1950V they are 350microns apart.

- The problem with the control of the FDC chiller was fixed (could not read out the temperature and flow) by Hovanes and the FDC LV and HV were turned on. So far we keep the HV at 2000V, some channels need more time for conditioning.


- Discussing the noise in the CDC, we want first to look at the noise in the FDC and then decide if and how we want to improve it (Fernando).